Problem #2-M&M Bags: Jordan interested in knowing the actual weight for a bag of plain M&Ms. He recently purchased thirty bags of milk chocolate lolain M&Ms. דhe measured weights for the sample are Riven below 49.79 48.98 5040 49.16 47,6 980 5023 5168 48,45 46.22 $043 43.80 46.94 4799 48.49 48.33 48.72 49.9 48.95 $1.71 5153 50.97 50.01 48.28 48.76.72 476 47.70 49.0 52.06 Define the objects in the population. Identify the variable and its units. State the sample size. B. Campute the mean and mesan for the weights C. Describe the shape of the distribution of weip ts in the histoaram below, D. Is the mean an adequate measure of a tyaical weight for this data set? EXPLAIN. Problem -Test Scores: Ernest and Emily both took Intro Stat l last spring, but with different professors. On their tespective Final exams Enest scored &9, while Emly scored 84 and Ernest is telling everyone that he scored higher than Emily. But Emily thinks that the sests were scored diflerently by the two prafessors, so she asked therm to provide the mean and standard deviation far the ir casies A. How many points abovehis dass average did Ernest store? Hew mar points above her dass average did Emily score? 8. Compute Ernests z score, then compute Emlys. C. Who has the more impressive test score (Ernest o Em? EXPLAIN Problem #3-Toothpaste Purchasers: Market researchers recently studied the age of the person in a househald wha is primarily responsible for purchasing toothpaste. es for Purchasers of Crest Brand 34 35 23 4452 45 2 244 33 55 3 28 48 28 34 33 245 31 60 54 3 S9 41 32 349 50 5 52 52 66 35 25 4 29 9 38 for Purchasers of Other Brands 9 61 2 A. Define the objects in each population Identify the varieble and its units. B. Enter the Ages for Purchasers of Crest Brand toothpaste into a list in your calculator for other technologv). Then compute the standard deviation and interquartile rangeR C. Enter the Ages for Purchasers of Other Brands af toothpaste into a list in your calculator (or other technology) Then campute the standard deviation and interquartile range(R D. Compare the standard deviatians AND compare the interquartile range of Ages between Crest and the Other brand (use comparison words, not just numbers E. In which group are the indvidual ages mare similar? Crest brand or other brands?) EXPLAIN.
Answer #1

ques 2

a) objects in the population are bags of plain M&Ms.
variable is weight of a bag of plain M&Ms
unit of variable is grams
sample size is size the sample used for measurement.here sample size is 30

b)

sorted data

X
46.22
46.72
46.94
47.61
47.67
47.7
47.98
48.28
48.33
48.45
48.49
48.72
48.74
48.95
48.98
49.16
49.4
49.69
49.79
49.8
49.8
50.01
50.23
50.4
50.43
50.97
51.53
51.68
51.71
52.06


mean =    ΣX/n = 1476.44/30= 49.21467

Median=0.5(n+1)th value =    15.5th value
=   49.07  

C) as we see from graph,shape of histogram is approx. normal.

d) yes,mean is adquate because graph is approx. normal.